# Increasing Batch Size Improves Convergence of Stochastic Gradient Descent with Momentum

### v1

@@ A.4 Training ResNet-18 on Tiny ImageNet
(a) Learning rate and batch size versus epochs
(b) Full gradient norm versus epochs
(c) Empirical loss versus epochs
(d) Test accuracy score versus epochs
Figure 4: (a) Schedulers for each optimizer with constant learning rates and constant batch size, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on Tiny ImageNet dataset.
(a) Learning rate and batch size versus epochs
(b) Full gradient norm versus epochs
(c) Empirical loss versus epochs
(d) Test accuracy score versus epochs
Figure 5: (a) Schedulers for each optimizer with constant learning rates and doubly increasing batch size every 25 epochs, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on Tiny ImageNet dataset.
(a) Learning rate and batch size versus epochs
(b) Full gradient norm versus epochs
(c) Empirical loss versus epochs
(d) Test accuracy score versus epochs
Figure 6: (a) Schedulers for each optimizer with constant learning rates and quadrupling increasing batch size every 50 epochs, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on Tiny ImageNet dataset.

### v2

@@ A.4 Training ResNet-18 on Tiny ImageNet
Figure 16: (a) Schedules for each optimizer with constant learning rates and constant batch size, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on the Tiny ImageNet dataset.
Figure 20: (a) Schedules for each optimizer with constant learning rates and batch size doubling every 25 epochs, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on the Tiny ImageNet dataset.
Figure 24: (a) Schedules for each optimizer with constant learning rates and batch size quadrupling every 50 epochs, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on the Tiny ImageNet dataset.
A.5 Visualization of SFO complexity (Tables 2-4)
(a) (a) SFO complexity to reach gradient norm thresholds with $b=8$ (corresponding to Table 2), (b) SFO complexity to reach gradient norm thresholds with $b=128$ (corresponding to Table 3), (c) SFO complexity to reach 70% test accuracy (corresponding to Table 4). Results are obtained using ResNet-18 on the CIFAR-100 dataset.
A.6 Results on wall-clock time
Since wall-clock time is an important practical metric of optimization efficiency, we additionally measured it under the same experimental settings as in Table 2, using CIFAR-100 with ResNet-18 and NSHB, with the maximum batch size set to 1024.
For the fixed batch size $b=8$ and the increasing batch size starting from $b=8$ and doubling every 20 epochs, the results are as follows:
Table 5: Wall-clock time to reach gradient norm thresholds ($b=8$)
Method Time to reach $\|\nabla f(\bm{\theta}_{t})\|<0.1$ (h:$m$:$s$) Time to reach $\|\nabla f(\bm{\theta}_{t})\|<0.05$ (h:$m$:$s$)
Fixed batch size ($b=8$) 2:22:58 3:40:05
Increasing batch size (initial $b=8$) 0:49:50 0:54:03
All experiments were conducted on the same computing server, using identical hardware and software environments to ensure a fair comparison. These improvements in wall-clock time are consistent with the reductions in SFO complexity reported in the main paper.
A.7 Assessing the validity of experimental learning rates in light of Theorem 3.2
The learning rate condition in Theorem 3.2 can be expressed as
Substituting $\beta=0.9$ into the first term yields approximately
$\eta\leq\frac{0.0270}{L}.$
This upper bound depends on the smoothness constant $L$, which is typically unknown and difficult to estimate accurately in practice. When $L$ is large, the bound becomes more restrictive, requiring a smaller $\eta$. Conversely, when $L$ is small, the condition allows for a relatively larger $\eta$, under which our empirical choice of $\eta=0.1$ still appears to fall within a theoretically reasonable range. Consequently, the learning rate setting adopted in our experiments can be regarded as theoretically justified across a wide range of possible values for $L$, and it is consistent with common empirical practices. Although the exact satisfaction of the theoretical bound cannot be rigorously verified, the consistent convergence observed in our experiments suggests that our setting lies within a reasonable range.